OpenStudy (anonymous):

how do i solve this equation please cos(t)^2=1/4 ????

OpenStudy (e.mccormick):

So, \(\cos^2(t)=1/4\)?

OpenStudy (e.mccormick):

If you think about it, anything squareed is positive, so the \(\pm\) square root issue comes into play. \[\cos^2t=\frac{1}{4} \implies \cos t =\pm \sqrt{\frac{1}{4}} \implies \cos t =\pm \frac{1}{2}\]

OpenStudy (e.mccormick):

So where on the unit circle is cosine 1/2 or -1/2?

OpenStudy (e.mccormick):

ANd did they give you a range for the answer or do you need all possible answers? Range means you list a few. All possible means you need to make an equation.

OpenStudy (e.mccormick):

I am off to class, but here is one last big clue:  The points on the unit circle you need.

OpenStudy (anonymous):

thank u!